OpenStudy (anonymous):

8/19 turned into a percentage

OpenStudy (dumbcow):

simply divide, then move decimal over 2 places 8/19 = .42105 --> 42.1%

OpenStudy (anonymous):

I got 42%

OpenStudy (dumbcow):

same thing depending on how you are rounding

OpenStudy (anonymous):

Nope no rounding I just got 42%

OpenStudy (dumbcow):

hmm no 42% = 42/100 = 21/50 this is not same as 8/19

OpenStudy (anonymous):

Read that again please.

OpenStudy (dumbcow):

??